# CDsampling: An R Package for Constrained D-Optimal Sampling in Paid Research Studies

### v2

@@ 3   Examples
Fisher Information $\mathbf{W}$ matrix diagonal elements of GLM W_func_GLM()
Fisher information matrix and determinant of GLM F_func_GLM()
Fdet_func_GLM()
Fisher information matrix of MLM at a specific design point Fi_func_MLM()
Fisher information matrix and determinant of MLM F_func_MLM()
Fdet_func_MLM()
Sampling Unconstrained D-optimal approximate allocation liftone_GLM()
liftone_MLM()
Constrained D-optimal approximate allocation liftone_constrained_GLM()
liftone_constrained_MLM()
Transfer from approximate allocation to exact allocation approxtoexact_constrained_func()
@@
Constrained uniform exact allocation bounded_uniform()
Fdet_unif()
Example Specific Find $I$ set for exact allocation conversion in examples iset_func_trauma()
iset_func_trial()
@@ 3.1   Applications in paid research study: $\tt trial\_data$ example
with $(\beta_{0},\beta_{1},\beta_{21},\beta_{22})=(0,3,3,3)$, where $i=1,\ldots,6$ stands for the $i$th covariate combination, and $j=1,\ldots,n_{i}$ is an index of patients who fall into the $i$th covariate combination. Figure 4 illustrates the distribution of treatment efficacy across different gender and age groups.
@@
In this example, it is posited that a sample of $n=200$ participants is desired from the population of $N=500$ volunteers due to budget constraints. The objective is to examine the variation in efficacy rates across gender and age demographics. Should a pilot study or analogous literature provide approximate values for the model parameters, a constrained lift-one algorithm may be employed to ascertain a locally D-optimal design. Conversely, if only partial parameter information is available, the expected values can be deduced from some prior distributions, and the constrained lift-one algorithm can be utilized to determine an EW D-optimal allocation (substituting $\mathbf{W}$ in the Fisher information matrix (1) with $E(\mathbf{W})$).
There are $m=6$ design points, corresponding to gender and age group combinations $(x_{gender\_i},$ $x_{age\_1i},x_{age\_2i})=(0,0,0)$, $(0,1,0)$, $(0,0,1)$, $(1,0,0)$, $(1,1,0)$, and $(1,0,1)$, respectively. The numbers of available volunteers in the six categories are $(N_{1},N_{2},\ldots,N_{6})=(50,40,10,$ 200,150,50)200,150,50)200 , 150 , 50 ). Our goal is to find the constrained D-optimal allocation $(w_{1},w_{2},\dots,w_{6})$ in the set of feasible allocations $S=\{(w_{1},\ldots,$ $w_{m})^{T}\in S_{0}\mid nw_{i}\leq N_{i},i=1,\ldots,m\}$.
@@
> beta = c(0, 3, 3, 3)> X=matrix(data=c(1,0,0,0,1,0,1,0,1,0,0,1,1,1,0,0,1,1,1,0,1,1,0,1), ncol=4, byrow=TRUE)> W=W_func_GLM(X=X, b=beta, link="logit")To define the number of design points, sample size, and constraints with $S$, we use the following R codes (see Section S3 in the Supplementary Materials of Huang et al. (2023) for details on finding $r_{i1}$ and $r_{i2}$):
@@
> set.seed(092)> approximate_design = liftone_constrained_GLM(X=X, W=W, g.con=g.con, g.dir=g.dir,+ g.rhs=g.rhs, lower.bound=lower.bound, upper.bound=upper.bound, reltol=1e-10,+ maxit=100, random=TRUE, nram=4, w00=NULL, epsilon=1e-8)The design output is presented below, with $\mathbf{w}$ representing the approximate D-optimal allocation:
@@
> exact_design = approxtoexact_constrained_func(n=200, w=approximate_design\$w, m=6,+ beta=beta, link=’logit’, X=X, Fdet_func=Fdet_func_GLM, iset_func=iset_func_trial)The sampling results are summarized below, with "allocation" representing the exact D-optimal allocation, which is the number of repetitions or subjects in each covariate combination, and "allocation.real" representing the corresponding proportion:
> print(exact_design)Optimal Sampling Results:================================================================================allocation :50.0, 40.0, 10.0, 100.0, 0.0, 0.0--------------------------------------------------------------------------------allocation.real :0.25, 0.2, 0.05, 0.5, 0.0, 0.0--------------------------------------------------------------------------------det.maximum :46.1012--------------------------------------------------------------------------------
@@
> set.seed(123)> approximate_design_EW = liftone_constrained_GLM(X=X, W=W.EW.unif,  g.con=g.con,+ g.dir=g.dir, g.rhs=g.rhs, lower.bound=lower.bound, upper.bound=upper.bound,+ reltol=1e-12, maxit=100, random=TRUE, nram=4, w00=NULL, epsilon=1e-10)> exact_design = approxtoexact_constrained_func(n=200,+ w=approximate_design_EW\$w, m=6, beta=beta, link=’logit’, X=X,+ Fdet_func=Fdet_func_GLM, iset_func=iset_func_trial)The output is summarized with print() function and presented below:
> print(exact_design_EW)Optimal Sampling Results:================================================================================allocation :48.0, 40.0, 10.0, 43.0, 19.0, 40.0--------------------------------------------------------------------------------allocation.real :0.2406, 0.2, 0.05, 0.2102, 0.0991, 0.2001--------------------------------------------------------------------------------det.maximum :25.59--------------------------------------------------------------------------------
In situations when the model parameters are unknown, the constrained uniform allocation is applicable. This method entails sampling an equal number of patients from each category within the given constraints. The selection criterion is $n_{i}=\min\{k,N_{i}\}$ or $\min\{k,N_{i}\}+1$ with $k$ satisfying $\sum_{i=1}^{m}\min\{k,N_{i}\}\leq n<\sum_{i=1}^{m}\min\{k+1,N_{i}\}$, where $N_{i}$ represents the maximum allowable number for each category. This is an example of a bounded design problem, where each category has an upper boundary. The function bounded_uniform() can be used to find the constrained uniform allocation with the trial_data example and "allocation" in the output representing the constrained uniform allocation.
> bounded_uniform(Ni=c(50, 40, 10, 200, 150, 50), nsample=200)Optimal Sampling Results:================================================================================allocation :38.0, 38.0, 10.0, 38.0, 38.0, 38.0--------------------------------------------------------------------------------
@@
> w00 = rep(1/200, 6)> unif_design = approxtoexact_constrained_func(n=200, w=w00, m=6, beta=NULL,+ link=NULL, X=NULL, Fdet_func=Fdet_func_unif, iset_func=iset_func_trial)> print(unif_design)Optimal Sampling Results:================================================================================allocation :38.0, 38.0, 10.0, 38.0, 38.0, 38.0--------------------------------------------------------------------------------allocation.real :0.005, 0.005, 0.005, 0.005, 0.005, 0.005--------------------------------------------------------------------------------det.maximum :792351680.0--------------------------------------------------------------------------------
The iset_func_trial() function in the CDsampling package is specifically designed for the trial_data example, which defines the set $I$ in step $1^{\circ}$ and step $2.4$ in the constrained approximate to exact algorithm depicted by Figure 3. This function can be replaced with customized functions according to users’ needs.
To perform a comparison analysis on different sampling strategies, including the constrained D-optimal allocation, the constrained EW D-optimal allocation with uniform priors, the constrained uniform allocation, simple random sample without replacement (SRSWOR), as well as the full data (all the $500$ patients), we simulate their responses $Y_{ij}$’s based on model (3) with parameter values $(0,3,3,3)$. We apply each sampling strategy to obtain $n=200$ observations out of $500$, and estimate the parameters using the $200$ observations. We use the root mean square error (RMSE) to measure the accuracy of the estimates (see Section 4 of Huang et al. (2023) for more theoretical and technical details). We repeat the procedure $100$ times and display the corresponding RMSEs in Figure 5 and simulation codes in Supplementary Section S3. Obviously, if we use the full data (all $500$ patients) to fit the model, its RMSE attains the lowest. Besides that, the constrained locally D-optimal allocation and the constrained EW D-optimal allocation have a little higher RMSEs than the full data estimates, but outperform SRSWOR and the constrained uniform allocation. The sampling strategy based on the constrained uniform allocation doesn’t need any model information and is a more robust sampling scheme, which is still better than SRSWOR.
@@ 3.2   Applications in paid research study: $\tt trauma\_data$ example
In the CDsampling package, trauma_data is a dataset of $N=802$ trauma patients from Chuang-Stein and Agresti (1997), stratified according to the trauma severity at the entry time of the study with $392$ mild and $410$ moderate/severe patients enrolled. The study involved four treatment groups determined by the dose level, $x_{i2}=1$ (Placebo), $2$ (Low dose), $3$ (Medium dose), and $4$ (High dose). Combining with severity grade ($x_{i1}=0$ for mild or $1$ for moderate/severe), there are $m=8$ distinct experimental settings with $(x_{i1},x_{i2})=(0,1),(0,2),(0,3),(0,4),(1,1),(1,2),(1,3),(1,4)$, respectively. The responses belong to five ordered categories, Death ($1$), Vegetative state ($2$), Major disability ($3$), Minor disability ($4$) and Good recovery ($5$), known as the Glasgow Outcome Scale (Jennett and Bond, 1975). Figure 6 shows the distribution of outcomes over severity grades and dose levels.
@@
In this example, we have $m=8$ design points, which are combinations of the two covariates: dose levels and severity grades. We aim to enroll $n=600$ patients from the $802$ available patients. The collection of feasible allocations is inherently constrained by the number of patients in different severity grades, defined as $S={(w_{1},\ldots,w_{8})^{\top}\in S_{0}\mid n(w_{1}+w_{2}+w_{3}+w_{4})\leq 392,n(w_{5}+w_{6}+w_{7}+w_{8})\leq 410}$.
The parameters fitted from the trauma_data are $\boldsymbol{\beta}=(\hat{\beta}_{11},\hat{\beta}_{21},\hat{\beta}_{31},$ $\hat{\beta}_{41},\hat{\beta}_{12},\hat{\beta}_{22},\hat{\beta}_{32},\hat{\beta}_{42},\hat{\beta}_{13},\hat{\beta}_{23},$ $\hat{\beta}_{33},\hat{\beta}_{43})^{\top}$ $=$ (4.047,2.225,0.302,1.386,4.214,3.519,2.420,(-4.047,-2.225,-0.302,1.386,4.214,3.519,2.420,( - 4.047 , - 2.225 , - 0.302 , 1.386 , 4.214 , 3.519 , 2.420 , $1.284,$ $-0.131,$ $-0.376,$ $-0.237,$ $-0.120)^{\top}$. The model can be written as the following four equations:
@@
> J=5; p=12; m=8;> beta = c(-4.047, -0.131, 4.214, -2.225, -0.376, 3.519, -0.302, -0.237,  2.420, 1.386,+ -0.120,  1.284)> Xi=rep(0,J*p*m); dim(Xi)=c(J,p,m);> Xi[,,1] = rbind(c( 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 1, 1, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 1, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,2] = rbind(c( 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 1, 2, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 2, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,3] = rbind(c( 1, 3, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 1, 3, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 3, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,4] = rbind(c( 1, 4, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 1, 4, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 4, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,5] = rbind(c( 1, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 1, 1, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 1, 1, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,6] = rbind(c( 1, 2, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 1, 2, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 2, 1, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,7] = rbind(c( 1, 3, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 1, 3, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 3, 1, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,8] = rbind(c( 1, 4, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 1, 4, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 4, 1, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))
@@
To define the sample size, the constraints, and the functions of lower and upper boundaries $r_{i1}$ and $r_{i2}$, we may use the following R codes (see Section S3 in the Supplementary Materials of Huang et al. (2023) for details on finding $r_{i1}$ and $r_{i2}$):

### v3

@@ 3   Examples
Model Calculating $\mathbf{W}$ matrix diagonal elements of generalized linear model (see Section 2.2); providing input for function liftone_GLM() and liftone_constrained_GLM(). W_func_GLM()
Calculating Fisher information matrix and its determinant of generalized linear model (see Example S1.1 in Supplementary Material). F_func_GLM(), Fdet_func_GLM()
Calculating Fisher information matrix of multinomial logit model at a specific design point (see Section 2.3); using as input of liftone_MLM() and liftone_constrained_MLM(). Fi_func_MLM()
Calculating Fisher information matrix and its determinant of multinomial logit model (see Example S1.2 in Supplementary Material). F_func_MLM(), Fdet_func_MLM()
Using in approxtoexact_constrained_func() to find constrained uniform exact allocation. Fdet_func_unif()
Sampling Finding unconstrained D-optimal approximate allocation for generalized linear model and multinomial logit model (see Section S2 in Supplementary Material). liftone_GLM(), liftone_MLM()
Finding constrained D-optimal approximate allocation for generalized linear model and multinomial logit model (see Section 3). liftone_constrained_GLM(), liftone_constrained_MLM()
Transferring approximate allocation to exact allocation (see Section 3). approxtoexact_constrained_func()
@@
Finding constrained uniform exact allocation for bounded constraint (see Section 3.1). bounded_uniform()
Example Specific Finding $I$ set for exact allocation conversion in trial_data and trauma_data examples (see Section 3 and Section S4). iset_func_trauma(), iset_func_trial()
@@ 3.1   Applications in paid research study: $\tt trial\_data$ example
with $(\beta_{0},\beta_{1},\beta_{21},\beta_{22})=(0,3,3,3)$, where $i=1,\ldots,6$ stands for the $i$th covariate combination, and $j=1,\ldots,n_{i}$ is an index of patients who fall into the $i$th covariate combination or sampling subgroups. Figure 4 illustrates the distribution of treatment efficacy across different gender and age groups.
@@
In this example, it is posited that a sample of $n=200$ participants is desired from the population of $N=500$ volunteers due to budget constraints. The objective is to examine the variation in efficacy rates across gender and age demographics. Should a pilot study or relevant literature provide approximate values for the model parameters, a constrained lift-one algorithm may be employed to find a locally D-optimal design. Conversely, if only partial parameter information is available, the expectation can be deduced from some prior distributions, and the constrained lift-one algorithm can be utilized to determine an EW D-optimal allocation (substituting $\mathbf{W}$ in the Fisher information matrix (1) with $E(\mathbf{W})$).
There are $m=6$ design points, corresponding to gender and age group combinations $(x_{gender\_i},$ $x_{age\_1i},x_{age\_2i})=(0,0,0)$, $(0,1,0)$, $(0,0,1)$, $(1,0,0)$, $(1,1,0)$, and $(1,0,1)$, respectively. The numbers of available volunteers in the six categories are $(N_{1},N_{2},\ldots,N_{6})=(50,40,10,$ 200,150,50)200,150,50)200 , 150 , 50 ). Our goal is to find the constrained D-optimal allocation $(w_{1},w_{2},\dots,w_{6})$ in the set of all feasible allocations $S=\{(w_{1},\ldots,$ $w_{m})^{T}\in S_{0}\mid nw_{i}\leq N_{i},i=1,\ldots,m\}$.
@@
> beta = c(0, 3, 3, 3) #coefficients> #design matrix> X=matrix(data=c(1,0,0,0,1,0,1,0,1,0,0,1,1,1,0,0,1,1,1,0,1,1,0,1), ncol=4, byrow=TRUE)> W=W_func_GLM(X=X, b=beta, link="logit") #find W as input of constrained liftoneTo define the number of design points, sample size, and constraints with $S$, we use the following R codes (see Section S3 in the Supplementary Material of Huang et al. (2023) for details on finding $r_{i1}$ and $r_{i2}$ in step 3 of the constrained liftone algorithm in Figure 2):
> rc = c(50, 40, 10, 200, 150, 50)/200 #available volunteers/sample size> m = 6 #design points> g.con = matrix(0,nrow=(2*m+1), ncol=m) #constraints> g.con[1,] = rep(1, m)> g.con[2:(m+1),] = diag(m)> g.con[(m+2):(2*m+1), ] = diag(m)> g.dir = c("==", rep("<=", m), rep(">=", m)) #direction> g.rhs = c(1, rc, rep(0, m)) #righ-hand side> #lower bound in step 3 of constrained liftone> lower.bound=function(i, w){+   rc = c(50, 40, 10, 200, 150, 50)/200+   m=length(w)+   temp = rep(0,m)+   temp[w>0]=1-pmin(1,rc[w>0])*(1-w[i])/w[w>0];+   temp[i]=0;+   max(0,temp);+   }> #upper bound in step 3 of constrained liftone> upper.bound=function(i, w){+   rc = c(50, 40, 10, 200, 150, 50)/200+   min(1,rc[i]);+   }
To identify the subgroups of the output D-optimal allocations, we may add an optional label for each of the $m=6$ covariatres combination or subgroups as “F, 18-25”, “F, 26-64”, “F, >=65”, “M, 18-2”, “M, 26-6”, “M, >=65” using the following codes:
> label = c("F, 18-25", "F, 26-64", "F, >=65", "M, 18-25", "M, 26-64", "M, >=65")
@@
> set.seed(092)> approximate_design = liftone_constrained_GLM(X=X, W=W, g.con=g.con, g.dir=g.dir,+ g.rhs=g.rhs, lower.bound=lower.bound, upper.bound=upper.bound, reltol=1e-10,+ maxit=100, random=TRUE, nram=4, w00=NULL, epsilon=1e-8, label=label)The design output is presented below:
> print(approximate_design)Optimal Sampling Results:================================================================================Optimal approximate allocation:   F, 18-25 F, 26-64 F, >=65 M, 18-25 M, 26-64 M, >=65w  0.25     0.2      0.05    0.5      0.0      0.0w0 0.25     0.0      0.05    0.7      0.0      0.0--------------------------------------------------------------------------------maximum :2.8813e-08--------------------------------------------------------------------------------convergence :TRUE--------------------------------------------------------------------------------itmax :9.0--------------------------------------------------------------------------------deriv.ans :0.0, 3.6017e-08, 4.8528e-07, -1.1525e-07, -1.0310e-07, -7.9507e-08--------------------------------------------------------------------------------gmax :0.0--------------------------------------------------------------------------------reason :"gmax <= 0"--------------------------------------------------------------------------------
The output includes several key components:
  • $\mathbf{w}$: reports the D-optimal approximate allocation.
  • $\mathbf{w}_{0}$: reports the random initial weights used to initialize the optimization.
  • maximum: reports the maximum determinant of Fisher information matrix.
  • reason: reports the termination criterion for the constrained lift-one algorithm including either “all derivative $\leq$ 0” or “gmax $\leq$ 0”, which corresponds to step $7^{\circ}$ and step $8^{\circ}$ in the constrained lift-one algorithm in Figure 2.
@@
> exact_design = approxtoexact_constrained_func(n=200, w=approximate_design\$w, m=6,+ beta=beta, link=’logit’, X=X, Fdet_func=Fdet_func_GLM, iset_func=iset_func_trial,+ label=label)> print(exact_design)Optimal Sampling Results:================================================================================Optimal exact allocation:                F, 18-25 F, 26-64 F, >=65 M, 18-25 M, 26-64 M, >=65allocation      50.0     40.0     10.0    100.0    0.0      0.0allocation.real 0.25     0.2      0.05    0.5      0.0      0.0--------------------------------------------------------------------------------det.maximum :46.1012--------------------------------------------------------------------------------
The output provides three key components for the sampling results:
  • allocation: reports the exact allocation of D-optimal sampling, specifying the number of subjects to sample from each subgroup.
  • allocation.real: reports the real-number approximate allocation used prior to integer conversion.
  • det.maximum: reports the maximum determinant of the Fisher information matrix by the optimal design.
In this example, the D-optimal exact allocation is to sample $50$ subjects from the “female, $18-25$” subgroup, $40$ subjects from the “female, $26-64$” subgroup, $10$ subjects from the “female, $\geq 65$” subgroup, and $100$ subjects from the “male, $18-25$” subgroup. Such a design may not explore all the design space, and lead to extreme design cases. In practice, allocating some subjects to the omitted subgroups “male, $26-64$” and “male, $\geq 65$” subgroups could improve the robustness and reduce the risk of overfitting.
@@
> set.seed(123)> approximate_design_EW = liftone_constrained_GLM(X=X, W=W.EW.unif,  g.con=g.con,+ g.dir=g.dir, g.rhs=g.rhs, lower.bound=lower.bound, upper.bound=upper.bound,+ reltol=1e-12, maxit=100, random=TRUE, nram=4, w00=NULL, epsilon=1e-10, label=label)> exact_design = approxtoexact_constrained_func(n=200,+ w=approximate_design_EW\$w, m=6, beta=beta, link=’logit’, X=X, Fdet_func=Fdet_func_GLM,+ iset_func=iset_func_trial, label=label)The output is summarized with print() function and presented below:
> print(exact_design_EW)Optimal Sampling Results:================================================================================Optimal exact allocation:                F, 18-25 F, 26-64 F, >=65 M, 18-25 M, 26-64 M, >=65allocation      48.0     40.0     10.0    43.0     19.0     40.0allocation.real 0.2406   0.2      0.05    0.2102   0.0991   0.2001--------------------------------------------------------------------------------det.maximum :25.59--------------------------------------------------------------------------------
In situations when the model parameters are unknown, the constrained uniform allocation is applicable. This method entails sampling an equal number of patients from each category within the given constraints. The selection criterion is $n_{i}=\min\{k,N_{i}\}$ or $\min\{k,N_{i}\}+1$ with $k$ satisfying $\sum_{i=1}^{m}\min\{k,N_{i}\}\leq n<\sum_{i=1}^{m}\min\{k+1,N_{i}\}$, where $N_{i}$ represents the maximum allowable number for each category. This is an example of a bounded design problem, where each category has an upper boundary. The function bounded_uniform() can be used to find the constrained uniform allocation with the trial_data example and “allocation” in the output representing the constrained uniform allocation.
> bounded_uniform(Ni=c(50, 40, 10, 200, 150, 50), nsample=200)Optimal Sampling Results:================================================================================Optimal exact allocation:           F, 18-25 F, 26-64 F, >=65 M, 18-25 M, 26-64 M, >=65allocation 38.0     38.0     10.0    38.0     38.0     38.0--------------------------------------------------------------------------------
@@
> w00 = rep(1/200, 6) #initial approximate allocation> unif_design = approxtoexact_constrained_func(n=200, w=w00, m=6, beta=NULL,+ link=NULL, X=NULL, Fdet_func=Fdet_func_unif, iset_func=iset_func_trial, label=label)> print(unif_design)Optimal Sampling Results:================================================================================Optimal exact allocation:                F, 18-25 F, 26-64 F, >=65 M, 18-25 M, 26-64 M, >=65allocation      38.0     38.0     10.0    38.0     38.0     38.0allocation.real 0.005    0.005    0.005   0.005    0.005    0.005--------------------------------------------------------------------------------det.maximum :792351680.0--------------------------------------------------------------------------------
[…]
To perform a comparison analysis on different sampling strategies, including the constrained D-optimal allocation, the constrained EW D-optimal allocation with uniform priors, the constrained uniform allocation, simple random sample without replacement (SRSWOR), as well as the full data (all the $500$ patients enrolled), we simulate their responses $Y_{ij}$’s based on model (3) with parameter values $(0,3,3,3)$. We apply each sampling strategy to obtain a sample of $n=200$ observations out of $500$, and estimate the parameters using the $200$ observations. The exception is the full data method, where estimation is performed using all $500$ patients. We use the root mean square error (RMSE) to measure the accuracy of the estimates (see Section 4 of Huang et al. (2023) for more theoretical and technical details). We repeat the procedure $100$ times and display the corresponding RMSEs in Figure 5 and simulation codes in Supplementary Material Section S3. Obviously, if we use the full data (all $500$ patients) to fit the model, its RMSE attains the lowest. Besides that, the constrained locally D-optimal allocation and the constrained EW D-optimal allocation have a little higher RMSEs than the full data estimates, but outperform SRSWOR and the constrained uniform allocation. The sampling strategy based on the constrained uniform allocation doesn’t need any model information and is a more robust sampling scheme, which is still better than SRSWOR.
@@ 3.2   Applications in paid research study: $\tt trauma\_data$ example
In the CDsampling package, trauma_data is a dataset of $N=802$ trauma patients from Chuang-Stein and Agresti (1997), stratified according to the trauma severity at the entry time of the study with $392$ mild and $410$ moderate/severe patients enrolled. The study involved four treatment groups determined by the dose level, $x_{i1}=1$ (Placebo), $2$ (Low dose), $3$ (Medium dose), and $4$ (High dose). Combining with severity grade ($x_{i2}=0$ for mild or $1$ for moderate/severe), there are $m=8$ distinct experimental settings with $(x_{i1},x_{i2})=(1,0),(2,0),(3,0),(4,0),(1,1),(2,1),(3,1),(4,1)$, respectively. The responses belong to five ordered categories, Death ($1$), Vegetative state ($2$), Major disability ($3$), Minor disability ($4$) and Good recovery ($5$), known as the Glasgow Outcome Scale (Jennett and Bond, 1975). Figure 6 shows the distribution of outcomes over severity grades and dose levels.
@@
In this example, we have $m=8$ subgroups, which are combinations of the two covariates categories: dose levels and severity grades. We aim to enroll $n=600$ patients from the $802$ available patients. The collection of feasible allocations is inherently constrained by the number of patients in different severity grades, defined as $S={(w_{1},\ldots,w_{8})^{\top}\in S_{0}\mid n(w_{1}+w_{2}+w_{3}+w_{4})\leq 392,n(w_{5}+w_{6}+w_{7}+w_{8})\leq 410}$. The constraints specify that in the sample, the number of patients with mild symptoms must not exceed $392$ across all dose levels, while those with moderate/severe symptoms must not exceed $410$.
The parameters fitted from the trauma_data are $\boldsymbol{\beta}=(\hat{\beta}_{11},\hat{\beta}_{12},\hat{\beta}_{13},\hat{\beta}_{21},\hat{\beta}_{22},\hat{\beta}_{23},\hat{\beta}_{31},\hat{\beta}_{32},\hat{\beta}_{33},\hat{\beta}_{41},$ $\hat{\beta}_{42},\hat{\beta}_{43})^{\top}=(-4.047,-0.131,4.214,-2.225,-0.376,3.519,-0.302,-0.237,2.420,1.386,-0.120,1.284)^{\top}$. The model can be written in the following format:
@@
> J=5; p=12; m=8; #response levels; parameters; subgroups> #coefficients> beta = c(-4.047, -0.131, 4.214, -2.225, -0.376, 3.519, -0.302, -0.237,  2.420, 1.386,+ -0.120,  1.284)> #define design matrix of 8 subgroups> Xi=rep(0,J*p*m); dim(Xi)=c(J,p,m);> Xi[,,1] = rbind(c( 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 1, 0, 0, 0, 0, 0, 0, 0),+                 c( 0, 0, 0, 0, 0, 0, 1, 1, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 0),+                 c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,2] = rbind(c( 1, 2, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 2, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 2, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,3] = rbind(c( 1, 3, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 3, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 3, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,4] = rbind(c( 1, 4, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 4, 0, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 4, 0, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 0),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,5] = rbind(c( 1, 1, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 1, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 1, 1, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 1, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,6] = rbind(c( 1, 2, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 2, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 2, 1, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 2, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,7] = rbind(c( 1, 3, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 3, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 3, 1, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 3, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))> Xi[,,8] = rbind(c( 1, 4, 1, 0, 0, 0, 0, 0, 0, 0, 0, 0), c( 0, 0, 0, 1, 4, 1, 0, 0, 0, 0, 0, 0),+                  c( 0, 0, 0, 0, 0, 0, 1, 4, 1, 0, 0, 0), c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 1, 4, 1),+                  c( 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0, 0))
@@
To define the sample size, the constraints, and the functions of lower and upper boundaries $r_{i1}$ and $r_{i2}$, we may use the following R codes (see Section S3 in the Supplementary Material of Huang et al. (2023) for details on finding $r_{i1}$ and $r_{i2}$):